# How Much Content Do LLMs Generate That Induces Cognitive Bias in Users?

### v1

@@ Abstract
Large language models (LLMs) are increasingly integrated into applications ranging from review summarization to medical diagnosis support, where they affect human decisions. Even though LLMs perform well in many tasks, they may also inherit societal or cognitive biases, which can inadvertently transfer to humans. We investigate when and how LLMs expose users to biased content and quantify its severity. Specifically, we assess three LLM families in summarization and news fact-checking tasks, evaluating how much LLMs stay consistent with their context and/or hallucinate. Our findings show that LLMs expose users to content that changes the sentiment of the context in 21.86% of the cases, hallucinates on post-knowledge-cutoff data questions in 57.33% of the cases, and primacy bias in 5.94% of the cases. We evaluate 18 distinct mitigation methods across three LLM families and find that targeted interventions can be effective. Given the prevalent use of LLMs in high-stakes domains, such as healthcare or legal analysis, our results highlight the need for robust technical safeguards and for developing user-centered interventions that address LLM limitations.
How Much Content Do LLMs Generate That Induces Cognitive Bias in Users?
Abeer Alessa, Akshaya Lakshminarasimhan, Param Somane, Julian Skirzynski, Julian McAuley, Jessica Echterhoff UC San Diego {aalessa, athenkarailakshminar, psomane, jskirzynski, jmcauley, jechterh}@ucsd.edu
@@ 1 Introduction
LLMs perform well across numerous tasks Albrecht et al. (2022) like content summarization Laban et al. (2023) , translation Elshin et al. (2024) , question-answering Lin et al. (2025), and sentiment analysis Zhang et al. (2024) including tasks with extremely large context Beltagy et al. (2020); Press et al. (2021); Touvron et al. (2023); Chen et al. (2023); Peng et al. (2023). In many of those tasks, humans rely on LLMs for daily decision-making support Rastogi et al. (2023); Li et al. (2022) as well as in expert contexts such as for writing policy documentsChoi et al. (2024) or summarizing medical documents (Spotnitz et al., 2024). However, these models have been shown to exhibit several societal biases Zhao et al. (2018); Nadeem et al. (2020); Liang et al. (2021); He et al. (2021), e.g., models favor specific genders Zhao et al. (2018), ethnicities Caliskan et al. (2017), or fall into similar decision-making-patterns as cognitively biased humans Echterhoff et al. (2024). Our specific contributions are the following
Figure 1: Our work evaluates how LLMs alter the source text, e.g., by changing its sentiment, to subsequently transfer the introduced bias to humans and affect their decision-making. Specifically, we evaluate how LLMs filter source content which leads to exposure bias for the user, how LLMs reframe the text which lead to human framing bias and how LLMs hallucinate which leads to authority/confirmation bias.
Figure 2: When LLMs process content for users to consume, they may change this content, e.g., by changing its sentiment or omitting some relevant parts. Exposure to the altered content may elicit cognitive biases such as positive framing bias or primacy bias, and subsequently have humans make different decisions than they would take if they saw the original text.
@@
    We analyze the character and extent of content alteration LLMs introduce in summarization and question-answering tasks. We find that LLMs change the framing of the text in roughly one of every five summaries and overly focus on the beginning of a source text once every 20 summaries. We also find that LLMs hallucinate in question-answering tasks in over half of the cases.
@@
    We propose specific metrics to capture the emergence of these content changes, such as a primacy bias score that captures the relative focus of the summary at the beginning of the source.
@@
    We provide a dataset of almost 3000 instances to analyze model hallucination past a model’s knowledge cutoff.  [^1] We publish this data at https://huggingface.co/datasets/sparklessszzz/NewsLensSync.
@@
    We evaluate 18 distinct mitigation methods and discover effective methods that target each specific type of content alteration.
Our goal is to highlight LLMs’ unreliability in various tasks and stress the importance of mitigating their errors, as they can negatively influence human users. We hope that our work will serve as a stepping stone for creating methods that improve LLMs and interventions that teach humans how to collaborate with LLMs thoughtfully.
@@ 2 Related Work
2.1 Social and Representational Biases in Language Models
Previous work shows the existence of different biases in LLMs. For example, prior work explores ways to evaluate human-like cognitive biases Jones and Steinhardt (2022); Echterhoff et al. (2024). Primacy and recency bias in LLMs have been shown in question-answering settings, where the models tend to be more accurate when the answer is presented in earlier or later documents Liu et al. (2024); Peysakhovich and Lerer (2023); Wang et al. (2023); Zhang et al. (2023); Eicher and Irgolič (2024). When LLMs are presented with options (e.g., A/B/C/D), they are shown to be more sensitive to the order Zheng et al. (2023a). Other studies have demonstrated the existence of framing bias in LLMs Leto et al. (2024). In framing bias, alterations in the tone, e.g., from positive to negative, can change the model’s decision-making Jones and Steinhardt (2022); Echterhoff et al. (2024).
In contrast to prior work that evaluates biases inherent in model behavior, we quantitatively assess the extent to which LLM responses alter the underlying content and how much bias they emit through these modifications.
@@ 2.2 Cognitive Effects of LLM-Generated Content on Humans
LLMs generate content to help users with all kinds of different tasks, such as summarization Laban et al. (2023), coding Tong and Zhang (2024), and writing Spangher et al. (2024) for personal or professional Draxler et al. (2023) use. When a model is given a specific task, it can alter, omit, or add information to produce the final model output. This can have an impact on a user in several ways. Previous work shows that LLMs can influence users by reinforcing users’ existing opinions Sharma et al. (2024), which can increase users’ tendency towards confirmation bias. The style in which LLMs present advice can significantly influence user perceptions Wester et al. (2024).
Prior work shows that altered headlines can introduce a framing effect on people in collective decision-making Abels et al. (2024). When experts use LLMs, they accept their suggestions without much revision Choi et al. (2024)
Given this insight, our study evaluates framing changes that the model introduces when generating a summary from ground-truth content, which subsequently could bias a user.
@@ 3 Background
LLMs may amplify different cognitive biases when users interpret their outputs. We focus on four biases:
@@
Framing bias refers to changes in human decisions when seeing a problem presented differently. People make different choices based on the wording of the questions, even if the actual choices presented remain unchanged Tversky and Kahneman (1981). Based on this prior work, we assume that a change in framing impacts the user and is hence to be avoided, e.g., when comparing the framing of a ground-truth context and its summarization.
@@
Primacy and Recency Bias. In primacy bias, humans prioritize the information they encounter first, while in recency bias, humans prioritize information they encountered most recently Murphy et al. (2006). This can result in skewed decision-making regardless of relevance or accuracy, but simply by order of encountered information Glenberg et al. (1980). When LLMs fall prey to primacy/recency bias Echterhoff et al. (2024), they might filter or feature information to the user leading to exposure bias. We assume that systematically highlighting any part of the context, whether it appears first, middle, or last, can influence the human user and should therefore be avoided.
@@
Confirmation and Authority Bias. Humans favor information confirming their beliefs or expectations Nickerson (1998). We assume that when a user prompts the model with belief-consistent input, and the model hallucinates in response, this can reinforce the user’s prior beliefs and lead to biased or misinformed decisions. In authority bias, people are more likely to trust information from perceived experts or authoritative sources Milgram (1963); Cialdini (2007); Güner Gültekin (2024). Research on human-AI interaction suggests that similar dynamics apply when people interact with artificial systems perceived as experts, which can create challenges when verifying facts Glickman and Sharot (2025). For example, when a user asks an LLM to verify if/when a particular legislation was passed, a user with strong prior beliefs may exhibit confirmation bias, while others may default to authority bias.
4 Quantifying Content Alteration in LLM Responses
@@
Due to the effects of the aforementioned biases, we propose the following metrics to capture unintended content alterations introduced by LLMs.
4.1 Framing Effects in Model-Generated Content
@@
We test if a model changes the sentiment of its summary compared to the context. For example, a model shifts the summary sentiment to positive, whereas the ground truth context was neutral. To evaluate these changes, we classify the framing of context $f_{c}$ (positive, negative, or neutral) before a summarization task, as well as on the generated answer summary by the model $f_{s}$. We use GPT-4o-mini to classify the framing of a text and provide details of its accuracy in Appendix A. We consider a model to be consistent if $f_{c}$ equals $f_{s}$. We consider the Bias-Inducing (BI) score as the Framing-change fraction, and denote it by $\varphi_{\mathrm{frame}}$ as below
@@
4.2 Primacy Effects in Model-Generated Content
We assess whether the model emphasizes a particular part of the review. For example, a model focuses only on the beginning of an Amazon user review that begins with praise but ends with criticism. We split the original context into three equal parts and calculate the cosine similarity using the multilingual-E5 embedding Wang et al. (2024) between each part and the summary. We denote the similarities of the summary and the beginning as $s_{b_{i}}$, the middle as $s_{m_{i}}$, and the end as $s_{e_{i}}$ for each instance $i$. We define biased examples as those in which the similarity to the beginning exceeds the similarity to the middle by a threshold of at least $\alpha$. We define the Primacy Bias Score as:
@@ 4.2 Primacy Effects in Model-Generated Content
with $\alpha=0.05$ for this study. A higher score indicates that a greater proportion of summaries focus more on the beginning of the source content compared to the middle.
4.3 Hallucination Effects in Model-Generated Content
We quantify whether a model generates hallucinated content when asked to fact-check news that lies outside its knowledge cutoff. For example, a model may wrongly state that an event never occurred, even though it did, but only after the model’s training had been completed. We use pre- and post-knowledge-cutoff data and prompt a model to verify whether specific news is true or false. We test both the accuracy of true, real-world news (Actual News Accuracy) and its falsified (negated) counterparts (Falsified News Accuracy). We prompt the model to evaluate the factual accuracy of a given news description. We propose to evaluate the strict accuracy in identifying when both the true news and the falsified news were correctly predicted.
For strict accuracy, we evaluate paired instances of true news and its falsified counterparts:
@@ 4.3 Hallucination Effects in Model-Generated Content
where $P_{j}$ is the set of paired instances in dataset $D_{j}$, $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{f}$ are the model’s predictions for the true and falsified versions of the same news respectively, and $T$ is the time-horizon (pre- or post-model-knowledge-cutoff).
@@
[…]
A larger $\delta_{H}$ indicates a greater chance of generating incorrect information when responding to queries beyond the model’s knowledge cutoff.
5 Mitigating Bias-Inducing Patterns in LLM Responses
To mitigate biased patterns in LLM responses, we build on our quantification of content alterations introduced in Section 4. We introduce targeted interventions designed to ensure balanced content coverage, reduce misleading emphasis, and maintain accurate representation in model outputs.
5.1 General Purpose Mitigations
@@
Prior works suggest that simple awareness prompts can mitigate certain biases Mair et al. (2014); Echterhoff et al. (2024). For example, in the summarization setting, we use the following prompt:
@@
5.2 Primacy Effects Mitigation
We group interventions into four families: Prompt-level cues, chunk-based summarization, attention/order re-ranking, and decoding-time control. Each family targets the tendency of LLMs to over-weight early context (primacy) or flip the source sentiment. A compact reference of all the methods that we experimented with is given in Appendix C.
Stand-out methods.
Weighted Summaries is a chunk-based approach that ensures balanced representation from the source text. The algorithm first splits the document into three segments (beginning, middle, and end). It then generates a partial summary for each segment independently, constraining the output length for each part according to a pre-allocated token budget (33%, 34%, and 33% of the total desired summary length). These partial summaries are then combined to form the final output.
In the Mirostat Decoding approach, the model’s sampling randomness (temperature) is dynamically adjusted during generation. This ensures the output text maintains a consistent level of unpredictability, which can prevent the model from focusing too narrowly on initial sentences and encourage broader coverage of the source text. We use a target entropy of $\mu^{\star}\!=\!2.0$ in our runs.
5.3 Mitigation for Framing Effects
In the Weighted Token Decoding approach, we directly influence the model’s word choice during generation. At each step, we modify the output probabilities by adding a logarithmic weight to the logits of specific tokens before sampling, in particular we down-weight tokens associated with negative sentiment to reduce framing shifts, as detailed in Appendix C. Liu et al. (2021); Dathathri et al. (2020)
5.4 Mitigation for Hallucination
The mitigation strategies for hallucination are summarized in Appendix C.
Knowledge Boundary Awareness. We propose a novel approach that explicitly defines the model’s knowledge cutoff date, preventing the fabrication of information about events beyond training data.
Epistemic Tagging. This approach requires models to express confidence levels alongside factual assertions, which forces self-evaluation of fact-checking abilities (i.e., the model additionally respond with its confidence level high or low for each of its responses.) Lin et al. (2022).
All methods are prompt-based interventions. Prompts details are in Appendix E.

### v2

# Quantifying Cognitive Bias Induction in LLM-Generated Content

@@ Abstract
Large language models (LLMs) are integrated into applications like shopping reviews, summarization, or medical diagnosis support, where their use affects human decisions. We investigate the extent to which LLMs expose users to biased content and demonstrate its effect on human decision-making. We assess five LLM families in summarization and news fact-checking tasks, evaluating the consistency of LLMs with their context and their tendency to hallucinate on a new self-updating dataset. Our findings show that LLMs expose users to content that changes the context’s sentiment in 26.42% of cases (framing bias), hallucinate on 60.33% of post-knowledge-cutoff questions, and highlight context from earlier parts of the prompt (primacy bias) in 10.12% of cases, averaged across all tested models. We further find that humans are 32% more likely to purchase the same product after reading a summary of the review generated by an LLM rather than the original review. To address these issues, we evaluate 18 mitigation methods across three LLM families and find the effectiveness of targeted interventions.
Quantifying Cognitive Bias Induction in LLM-Generated Content
Abeer Alessa, Param Somane*, Akshaya Lakshminarasimhan* Julian Skirzynski, Julian McAuley, Jessica Echterhoff UC San Diego {aalessa, psomane, athenkarailakshminar, jskirzynski, jmcauley, jechterh}@ucsd.edu
11footnotetext: Equal contribution.
@@ 1 Introduction
LLMs perform well across numerous tasks Albrecht et al. (2022), such as content summarization Laban et al. (2023), translation Elshin et al. (2024), sentiment analysis Zhang et al. (2024). In many of these tasks, humans rely on LLMs for daily decision-making support Rastogi et al. (2023); Li et al. (2022) in expert contexts such as writing policy documents Choi et al. (2024) or summarizing medical documents (Spotnitz et al., 2024). However, models have been shown to inherit or exhibit several societal biases Zhao et al. (2018); Nadeem et al. (2020); Liang et al. (2021); He et al. (2021), e.g., favoring specific genders Zhao et al. (2018) and ethnicities Caliskan et al. (2017), or learn similar decision-making-patterns as cognitively biased humans Echterhoff et al. (2024).
Any bias within the model can affect how textual context is processed for a human task and introduce biased content to users. Our work specifically focuses on qualifying and quantifying the extent to which biased content is induced by LLMs to users, as visualized in Figure 1 with examples shown in Figure 2. We summarize our contributions as follows:
Figure 1: LLMs alter information from a source text when performing a task for the user (e.g., when an LLM summary has a different sentiment compared to the original text). This model behavior introduces biased content to humans and can hence affect their decision-making. We evaluate how LLMs highlight source content, leading to exposure/primacy bias for users, how LLMs reframe sentiments, leading to framing bias for humans, and how LLMs hallucinate, leading to authority/confirmation bias.
Figure 2: When LLMs process context for users to consume, they may change its content, e.g., by changing its sentiment or omitting some relevant parts. Exposure to altered content may elicit cognitive biases, such as positive framing bias or primacy bias, and subsequently lead humans to make different decisions than they would if they saw the original text.
@@
    We analyze the extent of content alteration that LLMs introduce in summarization and news fact-checking tasks. We find that LLMs alter the sentiment or framing of a text in 26.42% of summaries, disproportionately focus on the beginning of a source text in 10.12% of summaries, and hallucinate in 60.33% of post-knowledge-cutoff news summaries for real-world fact-checking questions, averaged across all tested models.
@@
    We find that human users are affected by biases introduced by LLMs. Participants in our human-subject study were 32% more likely to purchase a product when they read its positively-framed LLM summary than when they read the original review.
@@
    We provide a dataset that enables analysis of model hallucination after a model’s cut-off date, which is self-updating to remain relevant for future models and prevent leakage into training data. At the time of writing this paper, it has 8,554 data points [^1] https://huggingface.co/datasets/sparklessszzz/NewsLensSync .
@@
    We evaluate 18 distinct mitigation methods and highlight which ones were effective for specific types of content alterations.
@@ 2 Related Work
2.1 Social and Representation Biases in LLMs
Previous work has demonstrated the presence of various biases in LLMs. Primacy and recency bias in LLMs have been shown in question-answering settings, where models tend to be more accurate when the answer is presented earlier or later in the documents Liu et al. (2024); Peysakhovich and Lerer (2023); Wang et al. (2023); Zhang et al. (2023); Eicher and Irgolič (2024). When presented with options, LLMs have been shown to exhibit sensitivity to their order Zheng et al. (2023a), and framing Leto et al. (2024). In framing bias, alterations in the tone, e.g., from positive to negative, can change the model’s decisions Jones and Steinhardt (2022); Echterhoff et al. (2024). In contrast to prior work, we quantitatively assess the extent to which LLM responses are biased.
@@ 2.2 Cognitive Effects of LLM-Generated Content on Humans
When given a specific task, a model can alter, omit, or add information to produce the final output, which can affect users in several ways Wester et al. (2024); Choi et al. (2024). Previous work demonstrates that LLMs can influence users by reinforcing their existing opinions Sharma et al. (2024), increasing confirmation bias, or influencing political decision-making Fisher et al. (2024). The style in which LLMs present advice can significantly influence users’ perceptions of its usefulnessWester et al. (2024). Altered headlines can introduce a framing effect on individuals in collective decision-making Abels et al. (2024). Choi et al. (2024) found that even when experts use LLMs, they tend to adopt the suggestions with fewer revisions.
Our work quantifies the extent of bias the model generates in its output compared to its context by LLMs, and finds that these content changes influence human purchasing decisions in a human-subject study.
@@ 3 Background
Humans are prone to mental shortcuts that negatively influence rational decision-making, known as cognitive biases. These mental shortcuts are often amplified by simple changes of text, framing, or highlighting that could be introduced by an LLM. We focus on three cognitive biases in our evaluation setup, which we briefly introduce here.
@@
Framing bias refers to changes in human decisions based on how a problem is presented. Individuals make different choices based on the wording of questions, even when the underlying options remain unchanged Tversky and Kahneman (1981). Based on prior work, we assume that a change in framing affects users and should be avoided because it can influence their decisions, e.g., when comparing the framing of a factual context with its summarization.
@@
Primacy Bias. Primacy bias occurs when humans prioritize information encountered first. Murphy et al. (2006). This bias skews decision-making based solely on the order in which information is encountered Glenberg et al. (1980). When LLMs exhibit primacy bias Echterhoff et al. (2024), they may filter or highlight information, leading to exposure bias. We assume that systematically highlighting any part of the context, regardless of its position, can influence users and should be avoided.
@@
Confirmation and Authority Bias. Individuals favor information confirming their existing beliefs or expectations Nickerson (1998). We assume that if a user prompts a model with belief-consistent input and the model subsequently hallucinates, this can reinforce the user’s prior beliefs, leading to biased or misinformed decisions. Authority bias causes individuals to trust information from perceived experts or authoritative sources Milgram (1963); Cialdini (2007); Gültekin (2024). Research on human-AI interaction suggests that similar dynamics apply when interacting with artificial systems perceived as experts, which can create challenges when verifying facts Glickman and Sharot (2025). For example, if a user asks an LLM to verify whether a particular legislation was passed, a user with strong prior beliefs may exhibit confirmation bias, while others may exhibit authority bias.
4 Quantifying Content Alteration in LLMs Responses
@@
To capture unintended content alterations introduced by LLMs, we propose the following metrics, motivated by the aforementioned human cognitive biases. We focus on content alteration introduced through changes in content framing, content filtering/highlighting, and content truthfulness.
Framing Effects in Model-Generated Content
@@
We test whether a model changes the sentiment of its summary compared to the source context. For example, a model may shift the summary sentiment to positive, whereas the source context was neutral. To evaluate these changes, we classify the framing of the source context $f_{c}$ (positive, negative, or neutral) before summarization, as well as the framing of the generated summary $f_{s}$. [^2] We use GPT-o4-mini to classify the framing of the text, after evaluating different models for this task. Details are provided in Appendix A.1.
We consider a model consistent if $f_{c}$ equals $f_{s}$. We define the bias-inducing score as the framing-change fraction, denoted by $\varphi_{\mathrm{frame}}$ as follows:
@@
Primacy Effects in Model-Generated Content
We assess whether a model emphasizes a specific part of its context. For example, a model might focus solely on the beginning of an Amazon user review that starts with praise but concludes with criticism. We divide the source context into three equal segments and calculate the cosine similarity, using embeddings Wang et al. (2024), between each segment and the summary. We denote the similarities of the summary and the beginning as $s_{b_{i}}$, the middle as $s_{m_{i}}$, and the end as $s_{e_{i}}$ for each instance $i$. We define biased examples as those where the similarity to the beginning exceeds the similarity to the middle by at least a threshold $\alpha$. This comparison is designed to isolate the primacy effect (over-reliance on initial content) from potential recency effects (over-reliance on final content), which can be measured separately. We define this primacy bias score as:
@@
with $\alpha=0.05$ for this study. A higher score indicates that a greater proportion of summaries focus more on the beginning of the source content than on the middle.
Hallucination in Model-Generated Content
We quantify whether a model generates hallucinated content when asked to fact-check news content past its knowledge cutoff. For example, a model may incorrectly state that an event never occurred, even if it did occur after the model’s training was completed. We prompt a model to verify the truth of specific news items. We test the accuracy of true, real-world news (Actual News Accuracy) and its falsified (negated) counterparts (Falsified News Accuracy). We evaluate the strict accuracy as correctly identifying both true and falsified news, with the paired instances of true news and its falsified counterparts:
@@
where $P_{j}$ is the set of paired instances in dataset $D_{j}$, $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{f}$ are the model’s predictions for the true and falsified versions of the same news item, respectively. $T$ represents the time horizon (pre- or post-model-knowledge-cutoff).
@@
$\delta_{H}=|s_{H}^{\mathit{pre}}-s_{H}^{\mathit{post}}|$ (4)
A larger $\delta_{H}$ indicates a higher fraction of generated incorrect/hallucinated information.
Models
We evaluate six open- and closed-source language models from five families to ensure our findings are generalizable across architectures. Our selection includes small open-source models (Phi-3-mini-4k-Instruct, Llama-3.2-3B-Instruct, Qwen3-4B-Instruct), a medium-sized model (Llama-3-8B-Instruct), a large open-source model (Gemma-3-27B-IT), and a closed-source model (GPT-3.5-turbo). Details on all models, model selection per task and our hyperparameter settings are provided in Appendix A.2. Appendix Table A2 shows knowledge cutoff dates for each model.
5 Datasets
News Interviews
News summaries should cover all key information, regardless of where it appears, and remain neutral, matching the tone of the original content. We use MediaSum, a dialogue summarization dataset comprising media interviews from NPR and CNN Zhu et al. (2021), to measure framing and primacy effects. We randomly select 1,000 data points with $\leq$ 4,000 tokens each, ensuring that the full content fits within a model’s context window. The descriptive statistics of text lengths are provided in the Appendix in Table A3.
Product Reviews
Previous work suggests that changing the framing of the product influences consumer purchase decisions Wei et al. (2024). To explore this effect in the context of LLM outputs, we use the Amazon Reviews dataset  Ni et al. (2019). This dataset comprises customer reviews from Amazon across various product categories, providing insights into product quality, usability, and customer satisfaction Ni et al. (2019). We sample 1,000 random examples from the Electronics category, ensuring token lengths of $\leq$ 4,000 to maintain full-context input for the model. Prior work notes that primacy bias is more observed in longer context windows Liu et al. (2024). Because most Amazon reviews are short, we evaluate the 1,000 longest reviews (by token length). Appendix Table A3 reports the descriptive statistics of the resulting dataset.
News Hallucination
We provide a novel dataset to quantify how models discern factual information across temporal boundaries and in both true and negated versions of news articles.
Pre-Knowledge Cutoff Data: To evaluate models on content before their knowledge cutoff, we sample 2,700 random data points from the “All the News” dataset all (2020), which includes articles from 27 major U.S. publications between 2016 and 2020.
Post-Knowledge Cutoff Data: To assess model performance on unseen events, we introduce a self-updating dataset, NewsLensSync, designed to evaluate hallucinations beyond model knowledge cutoffs. It is built using NewsAPI new , accessing live and historical content from 150,000 publishers globally. We restrict our queries to the keyword "politics," English-only articles, and dates after April 2025, being past all model knowledge cutoffs in our study.
Counterfactual versions are generated using a transformer model to falsify the descriptions of the new articles Anschütz et al. (2023). This results in paired authentic/falsified versions of the same news items which are manually validated for semantic consistency, producing matched authentic/negated pairs. Each data point includes a publication timestamp for filtering according to model cutoff dates. Further information on our self-updating dataset is described in Appendix A.3
6 Mitigation Strategies
We introduce targeted interventions designed to ensure balanced content coverage, reduce misleading emphasis, and maintain accuracy in model outputs. We analyze a total of 18 mitigation methods, but describe only the most promising in this section. All other mitigation methods are shown in Appendix A.6.
6.1 General Purpose Mitigation
@@
Simple awareness prompts can mitigate multiple biases Mair et al. (2014); Echterhoff et al. (2024). To use this prompt-based mitigation in our summarization setting, we use the following prompt:
@@
6.2 Primacy Effect Mitigation
We group interventions into four categories: prompt-level cues, chunk-based summarization, attention/order re-ranking, and decoding-time control. Each category targets the tendency of LLMs to overemphasize early context (primacy) or alter the source sentiment. All methods we experimented with are further described in Appendix A.6.
Weighted Summaries is a chunk-based approach that ensures balanced representation of the source text. The algorithm first splits the document into three segments (beginning, middle, and end). It then generates a partial summary for each segment independently, constraining the output length for each part according to a pre-allocated token budget (33%, 34%, and 33% of the total desired summary length). These partial summaries are then combined to form the final output.
In Mirostat Decoding, the model’s temperature is dynamically adjusted during generation via a feedback loop (formula in Appendix Table A5). This ensures the output text maintains a consistent level of unpredictability, which prevents the model from focusing too narrowly on initial sentences and encourages broader coverage of the source text. We use a target entropy of $\mu^{\star}\!=\!2.0$ in our runs, similar to Basu et al. (2021).
6.3 Framing Effect Mitigation
Weighted Token Decoding directly influences the model’s token selection during the generation process. At each step, we modify the output probabilities by adding a logarithmic weight to the logits of specific tokens before sampling Liu et al. (2021); Dathathri et al. (2020). In particular, we down-weight tokens associated with negative sentiment to reduce framing shifts, as detailed in Appendix A.6.
6.4 Hallucination Mitigation
We analyse Knowledge Boundary Awareness, which explicitly defines the model’s knowledge cutoff date in the context to establish clear temporal limits on what the model can reasonably know. Epistemic Tagging requires models to express confidence levels alongside factual assertions, forcing self-evaluation of fact-checking abilities (i.e., the model additionally responds with a high or low confidence level for each response) Lin et al. (2022).
Hallucination mitigation strategies are detailed in Appendix A.6 and prompt details provided in Appendix A.8.